### v3

@@ 1 Introduction
Generative AI and large language models (LLMs) have the potential to drastically improve the landscape of computing and programming education by powering next-generation educational technologies. This potential lies in the advanced capabilities of state-of-the-art models—like OpenAI’s GPT-4 [1] and ChatGPT (based on GPT-3.5[2]—to automatically generate high-quality personalized content and feedback for students [3, 4, 5]. A series of recent works have already shown us sparks of their capabilities for various programming education scenarios, including generating new programming assignments [6, 7], providing code explanations [8, 6], repairing buggy programs [9, 10], enhancing programming-error-messages [11, 10], and acting as pair programmer [12, 13]. In this paper, we investigate the role of LLMs in providing human tutor-style programming hints to help students resolve errors in their buggy programs. More concretely, given a programming task and a student’s buggy program, we want to generate natural language hints to help the student resolve bug(s) and make progress, inspired by how a human tutor would give pedagogical feedback. With the current scale of enrollments in introductory programming courses [14], it has become infeasible for human tutors to promptly provide individualized feedback to students, thereby motivating the need to develop automatic feedback generation techniques. To this end, we aim to leverage generative AI and LLMs for automating human tutor-style programming feedback to support students’ learning and reduce human tutors’ workload. Recent works have studied state-of-the-art LLMs for generating various forms of programming feedback for students, including detailed explanations about bugs or single-sentence hints [4, 11, 10]. Despite promising initial results, the overall quality of feedback generated by LLMs is substantially inferior to that of human tutors and not yet ready for deployment in real-life classroom settings. For instance, a recent benchmark study in [4] evaluated GPT-4 in generating hints for buggy programs on introductory Python programming tasks and assessed its quality performance using expert annotations – GPT-4’s performance in terms of hints quality is only about $60\%$ in contrast to human tutors’s performance of over $90\%$. This performance gap between GPT-4 vs. human tutors can be attributed to several factors, as discussed next. First, state-of-the-art models still struggle with symbolic reasoning and program execution abilities crucial for understanding the underlying bugs and possible student misconceptions [3, 4, 5, 15]. Second, these models also suffer from hallucination issues and the generated feedback text—even though seemingly plausible—may contain inaccurate information that could have detrimental effects on students’ learning [15, 16, 17]. Third, these models still lack a calibration mechanism to decide whether the generated content is of high quality or not [10]; in particular, they are unable to do a human tutor-style reasoning from a student’s perspective and judge if the generated feedback would likely help the student.
Given a string S, check if it is palindrome or not.
Expected Time Complexity: $O(|S|)$.
@@
1#User function Template for python3 2class Solution: 3  def reverse(self, b, e, S): 4    while b < e: 5      S[b], S[e] = S[e], S[b] 6      e = e - 1 7      b = b + 1 8    return S 9 10  def isPalindrome(self, S): 11    S = list(S) 12    beg = 0 13    end = len(S) - 1 14    rev = self.reverse(beg,end,S) 15    if S == rev: 16      return 1 17    return 0 18 19#{ Driver Code Starts 20#Initial Template for Python 3 21if __name__ == ’__main__’: 22  T=int(input()) 23  for i in range(T): 24    S = input() 25    ob = Solution() 26    answer = ob.isPalindrome(S) 27    print(answer) 28# } Driver Code Ends (b) Student’s buggy program 1#User function Template for python3 2class Solution: 3  def reverse(self, b, e, S): 4    while b < e: 5      S[b], S[e] = S[e], S[b] 6      e = e - 1 7      b = b + 1 8    return S 9 10  def isPalindrome(self, S): 11    S = list(S) 12    beg = 0 13    end = len(S) - 1 14    rev = self.reverse(beg,end,S.copy()) 15    if S == rev: 16      return 1 17    return 0 18 19#{ Driver Code Starts 20#Initial Template for Python 3 21if __name__ == ’__main__’: 22  T=int(input()) 23  for i in range(T): 24    S = input() 25    ob = Solution() 26    answer = ob.isPalindrome(S) 27    print(answer) 28# } Driver Code Ends (c) Fixed program For Input: hq
@@
The bug in the program is that the original string S is being modified in the reverse function. When you compare the reversed string with the original string, they are always the same because they are both pointing to the same modified list. The fix is to pass a copy of the original string to the reverse function so that the original string remains unmodified. (e) Detailed explanation Consider the effect of Python’s pass-by-reference behavior on mutable objects like lists when you’re reversing the string. (f) Single-sentence hint Refer to caption (g) Validation
@@ 1.1 Our Approach and Contributions
In this paper, we seek to push the limits of generative AI and state-of-the-art LLMs toward providing high-quality programming hints. Given a base model, this would require improving the model’s abilities at input-level by developing better prompting strategies [18], at output-level by developing mechanisms to validate the generated content [10, 19, 20], or at model-level itself by fine-tuning (when considering open-source models [21]). In our work, we consider OpenAI’s GPT-4 [1] as the base model—the latest model presumably with over a trillion parameters—as it has shown to drastically improve existing models across various programming education scenarios [4]. We develop a novel technique, GPT4Hints-GPT3.5Val, to provide human tutor-style high-quality programming hints. Our technique leverages the GPT-4 model in the role of a “tutor” to generate hints and boosts the generative quality at the input level by prompting it with symbolic information of failing test cases and fixed programs. At the output level, it further validates the hint quality by leveraging the GPT-3.5 model as a “student” to simulate the potential utility of providing this feedback to human students. This validation step is designed to provide a quality assurance layer and decides whether the generated feedback should be provided to the human student or not – thereby trading off coverage (how many students are given automatic feedback) and precision (quality of the given feedback). We show the efficacy of our technique by conducting an extensive evaluation using three real-world datasets of Python programs covering a variety of concepts ranging from writing basic algorithms to regular expressions and data analysis using pandas [22]. Figures 1 and 2 showcase GPT4Hints-GPT3.5Val on two different buggy programs. [^1] When presenting these illustrative examples in this paper, we slightly obfuscate the students’ buggy programs to avoid showing exact real-world programs. We do so by altering variable names and formatting conventions while preserving the original bugs exactly the same, as considered in related works [10, 4]. Accordingly, if needed, we apply the same adjustments to the generated output to maintain consistency with these alterations. More broadly, our work makes the following contributions in leveraging generative AI and LLMs for computing and programming education:
@@
The motivation of the problem is to investigate any evidence of a link between vaccine efficacy and sex of the child. For this, you should compute the ratio of the number of children who contracted chickenpox but were vaccinated against it (at least one varicella dose) versus those who were vaccinated but did not contract chicken pox. Return results by sex.
Design your function so that it returns a dictionary in the following format:
@@
1def chickenpox_by_sex(): 2  import pandas as pd 3  import numpy as np 4  df = pd.read_csv(’assets/data.csv’, index_col=0) 5  male = df[df[’SEX’] == 1] 6  vac_m = male[male[’P_NUMVRC’] == 1] 7  cpox_m = vac_m[vac_m[’HAD_CPOX’] == 1] 8  no_m = vac_m[vac_m[’HAD_CPOX’] == 2] 9  total_cpox_m = cpox_m[’SEX’].count() 10  total_nocpox_m = no_m[’SEX’].count() 11  m = total_cpox_m/total_nocpox_m 12 13   ... code omitted to save space ... 14 15  ans={"male":  m, 16       "female": f} 17 18  return ans (b) Student’s buggy program 1def chickenpox_by_sex(): 2  import pandas as pd 3  import numpy as np 4  df = pd.read_csv(’assets/data.csv’, index_col=0) 5  male = df[df[’SEX’] == 1] 6  vac_m = male[male[’P_NUMVRC’] >= 1] 7  cpox_m = vac_m[vac_m[’HAD_CPOX’] == 1] 8  no_m = vac_m[vac_m[’HAD_CPOX’] == 2] 9  total_cpox_m = cpox_m[’SEX’].count() 10  total_nocpox_m = no_m[’SEX’].count() 11  m = total_cpox_m/total_nocpox_m 12 13   ... code omitted to save space ... 14 15  ans={"male":  m, 16       "female": f} 17 18  return ans (c) Fixed program Your Code’s
@@
The bug in the program is in the line where you are filtering the male children who were vaccinated. You are only considering those who received exactly one dose of the vaccine (‘male[‘P_NUMVRC’] == 1’), but the problem statement asks for children who received at least one dose of the vaccine. The fix is to change the condition to ‘male[‘P_NUMVRC’] >= 1’. (e) Detailed explanation Check the conditions you are using to filter the vaccinated children, especially the number of doses they received. (f) Single-sentence hint Refer to caption (g) Validation
@@ 2 Problem Setup
Programming task and student’s buggy program as input. We start with a programming task $\mathcal{T}$ and a buggy program $\mathcal{P}_{\textnormal{b}}$. A task $\mathcal{T}$, such as shown in Figures  0(a) and  1(a), is represented by a textual description of the programming problem. Additionally, this description encompasses all requisite information essential for problem solving, such as expected algorithm complexity and any constraints on input, as applicable. In cases where the task necessitates interaction with an external file, $\mathcal{T}$ should also contain all pertinent information of that file crucial for solving the problem, such as the file’s format or structure. $\mathcal{P}_{\textnormal{b}}$, as illustrated in Figures  0(b) and  1(b), is an unsuccessful attempt of the student to solve $\mathcal{T}$. This program fails to pass at least one of the test cases in the test suite for $\mathcal{T}$. In general, $\mathcal{P}_{\textnormal{b}}$ may contain one or multiple errors, spanning various error types including syntax and semantic errors.
@@ 3.1 Stage-1: Generate Symbolic Data
Overview and intuition. As discussed in Section 1, there remains a notable performance gap between state-of-the-art generative AI models and human tutors regarding hint generation. One key factor contributing to this disparity is the inability to do symbolic reasoning and program execution. GPT-4 lacks the capability to execute the given code to retrieve an output, which could help it gain deeper understanding of the underlying bugs. To mitigate this gap, we employ external tools to execute programs and extract useful symbolic information. We then supply this information to GPT-4 for feedback generation. Our approach centers on leveraging two categories of symbolic data: failing test cases and fixed programs. Input/output for a failing test case. To highlight the error in the buggy program $\mathcal{P}_{\textnormal{b}}$, we provide GPT-4 with a test case for which $\mathcal{P}_{\textnormal{b}}$ fails to produce the expected output. To acquire this test case, we run $\mathcal{P}_{\textnormal{b}}$ on the existing test suite given for the corresponding task $\mathcal{T}$. The first test case in which $\mathcal{P}_{\textnormal{b}}$ fails is selected. We denote the triplet comprising this input, the output generated by $\mathcal{P}_{\textnormal{b}}$, and the expected output, as $\omega$ and include it in the prompt for feedback generation. Fixed program. The fixed program, denoted as $\mathcal{P}_{\textnormal{f}}$, is generated using GPT-4, employing a procedure adapted from the work in [10]. To be more specific, we initiate the process by requesting the model to produce $10$ independent fixed programs. For this purpose, we include $\mathcal{T}$ and $\mathcal{P}_{\textnormal{b}}$ in the prompt [^3] The prompt used here has the same format as shown in Figure 4 (third prompt). to ask for 10 outputs (each output contains a fixed program) with the hyperparameter temperature set to $0.5$. Then, from this set of $10$, we take the programs that pass the test suite for $\mathcal{T}$ and among them, identify $\mathcal{P}_{\textnormal{f}}$ as the one with the smallest token-edit distance w.r.t. $\mathcal{P}_{\textnormal{b}}$. To compute the token-edit distance between two programs, we first tokenize them using Pygments library [31] and then calculate the Levenshtein edit distance based on the tokenized strings. If $\mathcal{P}_{\textnormal{f}}$ is found, we include it in the prompt for feedback generation. If, however, none of the generated programs is correct, we opt to exclude this symbolic information from the prompt.
@@ 3.2 Stage-2: Generate Feedback
Overview and intuition. In this stage, we aim to obtain a human tutor-style hint $\mathcal{H}$ as feedback to be given to the student, as previously mentioned in Section 2. In addition to our request for a hint $\mathcal{H}$ from GPT-4, we also ask for a detailed explanation, denoted as $\mathcal{X}$, for the bugs in $\mathcal{P}_{\textnormal{b}}$. The reason to ask for this explanation draws inspiration from Chain-of-Thought [18], an established method renowned for enhancing the reasoning capabilities of LLMs. The essence of the Chain-of-Thought approach lies in encouraging LLMs to explain their thought process meticulously, step by step, prior to presenting the final output. Within the specific context of hint generation, we allow the model to elaborate its reasoning through $\mathcal{X}$ before coming up with the concise single-sentence hint $\mathcal{H}$, which is essentially an abstracted version of the explanation. Furthermore, $\mathcal{X}$ will also play a pivotal role in the subsequent feedback validation stage, which will be elaborated upon in Section 3.3. Prompt for feedback generation. In Figure 4 (first prompt), we provide our prompt for generating feedback. This prompt comprises the problem description for $\mathcal{T}$, the buggy program $\mathcal{P}_{\textnormal{b}}$, the symbolic information as extracted from the previous stage, and a request for an explanation $\mathcal{X}$ along with a hint $\mathcal{H}$. To get a response from GPT-4, we use this prompt while configuring the hyperparameter temperature to 00, indicating our preference for the most probable answer. All other hyperparameters are kept at their default settings. Following this, $\mathcal{X}$ and $\mathcal{H}$ are then extracted automatically from the output.

### v4

@@ 1 Introduction
Generative AI and large language models (LLMs) have the potential to drastically improve the landscape of computing and programming education by powering next-generation educational technologies. This potential lies in the advanced capabilities of state-of-the-art models—like OpenAI’s GPT-4 [1] and ChatGPT (based on GPT-3.5[2]—to automatically generate high-quality personalized content and feedback for students [3, 4, 5]. A series of recent works have already shown us sparks of their capabilities for various programming education scenarios, including generating new programming assignments [6, 7], providing code explanations [8, 6], repairing buggy programs [9, 10], enhancing programming-error-messages [11, 10], and acting as pair programmer [12, 13].
In this paper, we investigate the role of LLMs in providing human tutor-style programming hints to help students resolve errors in their buggy programs. More concretely, given a programming task and a student’s buggy program, we want to generate natural language hints to help the student resolve bug(s) and make progress, inspired by how a human tutor would give pedagogical feedback. With the current scale of enrollments in introductory programming courses [14], it has become infeasible for human tutors to promptly provide individualized feedback to students, thereby motivating the need to develop automatic feedback generation techniques. To this end, we aim to leverage generative AI and LLMs for automating human tutor-style programming feedback to support students’ learning and reduce human tutors’ workload.
@@
Recent works have studied state-of-the-art LLMs for generating various forms of programming feedback for students, including detailed explanations about bugs or single-sentence hints [4, 11, 10]. Despite promising initial results, the overall quality of feedback generated by LLMs is substantially inferior to that of human tutors and not yet ready for deployment in real-life classroom settings. For instance, a recent benchmark study in [4] evaluated GPT-4 in generating hints for buggy programs on introductory Python programming tasks and assessed its quality performance using expert annotations – GPT-4’s performance in terms of hints quality is only about $60\%$ in contrast to human tutors’s performance of over $90\%$. This performance gap between GPT-4 vs. human tutors can be attributed to several factors, as discussed next. First, state-of-the-art models still struggle with symbolic reasoning and program execution abilities crucial for understanding the underlying bugs and possible student misconceptions [3, 4, 5, 15]. Second, these models also suffer from hallucination issues and the generated feedback text—even though seemingly plausible—may contain inaccurate information that could have detrimental effects on students’ learning [15, 16, 17]. Third, these models still lack a calibration mechanism to decide whether the generated content is of high quality or not [10]; in particular, they are unable to do a human tutor-style reasoning from a student’s perspective and judge if the generated feedback would likely help the student.
Given a string S, check if it is palindrome or not. Expected Time Complexity: $O(|S|)$.
@@
1#User function Template for python3 2class Solution: 3 def reverse(self, b, e, S): 4 while b < e: 5 S[b], S[e] = S[e], S[b] 6 e = e - 1 7 b = b + 1 8 return S 9 10 def isPalindrome(self, S): 11 S = list(S) 12 beg = 0 13 end = len(S) - 1 14 rev = self.reverse(beg,end,S) 15 if S == rev: 16 return 1 17 return 0 18 19#{ Driver Code Starts 20#Initial Template for Python 3 21if __name__ == ’__main__’: 22 T=int(input()) 23 for i in range(T): 24 S = input() 25 ob = Solution() 26 answer = ob.isPalindrome(S) 27 print(answer) 28# } Driver Code Ends
(b) Student’s buggy program
1#User function Template for python3 2class Solution: 3 def reverse(self, b, e, S): 4 while b < e: 5 S[b], S[e] = S[e], S[b] 6 e = e - 1 7 b = b + 1 8 return S 9 10 def isPalindrome(self, S): 11 S = list(S) 12 beg = 0 13 end = len(S) - 1 14 rev = self.reverse(beg,end,S.copy()) 15 if S == rev: 16 return 1 17 return 0 18 19#{ Driver Code Starts 20#Initial Template for Python 3 21if __name__ == ’__main__’: 22 T=int(input()) 23 for i in range(T): 24 S = input() 25 ob = Solution() 26 answer = ob.isPalindrome(S) 27 print(answer) 28# } Driver Code Ends
(c) Fixed program
For Input: hq
@@
The bug in the program is that the original string S is being modified in the reverse function. When you compare the reversed string with the original string, they are always the same because they are both pointing to the same modified list. The fix is to pass a copy of the original string to the reverse function so that the original string remains unmodified.
(e) Detailed explanation
Consider the effect of Python’s pass-by-reference behavior on mutable objects like lists when you’re reversing the string.
(f) Single-sentence hint
(g) Validation
@@ 1.1 Our Approach and Contributions
In this paper, we seek to push the limits of generative AI and state-of-the-art LLMs toward providing high-quality programming hints. Given a base model, this would require improving the model’s abilities at input-level by developing better prompting strategies [18], at output-level by developing mechanisms to validate the generated content [10, 19, 20], or at model-level itself by fine-tuning (when considering open-source models [21]). In our work, we consider OpenAI’s GPT-4 [1] as the base model—the latest model presumably with over a trillion parameters—as it has shown to drastically improve existing models across various programming education scenarios [4].
We develop a novel technique, GPT4Hints-GPT3.5Val, to provide human tutor-style high-quality programming hints. Our technique leverages the GPT-4 model in the role of a “tutor” to generate hints and boosts the generative quality at the input level by prompting it with symbolic information of failing test cases and fixed programs. At the output level, it further validates the hint quality by leveraging the GPT-3.5 model as a “student” to simulate the potential utility of providing this feedback to human students. This validation step is designed to provide a quality assurance layer and decides whether the generated feedback should be provided to the human student or not – thereby trading off coverage (how many students are given automatic feedback) and precision (quality of the given feedback). We show the efficacy of our technique by conducting an extensive evaluation using three real-world datasets of Python programs covering a variety of concepts ranging from writing basic algorithms to regular expressions and data analysis using pandas [22]. Figures 1 and 2 showcase GPT4Hints-GPT3.5Val on two different buggy programs. [^1] When presenting these illustrative examples in this paper, we slightly obfuscate the students’ buggy programs to avoid showing exact real-world programs. We do so by altering variable names and formatting conventions while preserving the original bugs exactly the same, as considered in related works [10, 4]. Accordingly, if needed, we apply the same adjustments to the generated output to maintain consistency with these alterations. More broadly, our work makes the following contributions in leveraging generative AI and LLMs for computing and programming education:
@@
The motivation of the problem is to investigate any evidence of a link between vaccine efficacy and sex of the child. For this, you should compute the ratio of the number of children who contracted chickenpox but were vaccinated against it (at least one varicella dose) versus those who were vaccinated but did not contract chicken pox. Return results by sex. Design your function so that it returns a dictionary in the following format:
@@
1def chickenpox_by_sex(): 2 import pandas as pd 3 import numpy as np 4 df = pd.read_csv(’assets/data.csv’, index_col=0) 5 male = df[df[’SEX’] == 1] 6 vac_m = male[male[’P_NUMVRC’] == 1] 7 cpox_m = vac_m[vac_m[’HAD_CPOX’] == 1] 8 no_m = vac_m[vac_m[’HAD_CPOX’] == 2] 9 total_cpox_m = cpox_m[’SEX’].count() 10 total_nocpox_m = no_m[’SEX’].count() 11 m = total_cpox_m/total_nocpox_m 12 13 ... code omitted to save space ... 14 15 ans={"male": m, 16 "female": f} 17 18 return ans
(b) Student’s buggy program
1def chickenpox_by_sex(): 2 import pandas as pd 3 import numpy as np 4 df = pd.read_csv(’assets/data.csv’, index_col=0) 5 male = df[df[’SEX’] == 1] 6 vac_m = male[male[’P_NUMVRC’] >= 1] 7 cpox_m = vac_m[vac_m[’HAD_CPOX’] == 1] 8 no_m = vac_m[vac_m[’HAD_CPOX’] == 2] 9 total_cpox_m = cpox_m[’SEX’].count() 10 total_nocpox_m = no_m[’SEX’].count() 11 m = total_cpox_m/total_nocpox_m 12 13 ... code omitted to save space ... 14 15 ans={"male": m, 16 "female": f} 17 18 return ans
(c) Fixed program
Your Code’s
@@
The bug in the program is in the line where you are filtering the male children who were vaccinated. You are only considering those who received exactly one dose of the vaccine (‘male[‘P_NUMVRC’] == 1’), but the problem statement asks for children who received at least one dose of the vaccine. The fix is to change the condition to ‘male[‘P_NUMVRC’] >= 1’.
(e) Detailed explanation
Check the conditions you are using to filter the vaccinated children, especially the number of doses they received.
(f) Single-sentence hint
(g) Validation
@@ 2 Problem Setup
Programming task and student’s buggy program as input. We start with a programming task $\mathcal{T}$ and a buggy program $\mathcal{P}_{\textnormal{b}}$. A task $\mathcal{T}$, such as shown in Figures  1(a) and  2(a), is represented by a textual description of the programming problem. Additionally, this description encompasses all requisite information essential for problem solving, such as expected algorithm complexity and any constraints on input, as applicable. In cases where the task necessitates interaction with an external file, $\mathcal{T}$ should also contain all pertinent information of that file crucial for solving the problem, such as the file’s format or structure. $\mathcal{P}_{\textnormal{b}}$, as illustrated in Figures  1(b) and  2(b), is an unsuccessful attempt of the student to solve $\mathcal{T}$. This program fails to pass at least one of the test cases in the test suite for $\mathcal{T}$. In general, $\mathcal{P}_{\textnormal{b}}$ may contain one or multiple errors, spanning various error types including syntax and semantic errors.
@@ 3.1 Stage-1: Generate Symbolic Data
Overview and intuition. As discussed in Section 1, there remains a notable performance gap between state-of-the-art generative AI models and human tutors regarding hint generation. One key factor contributing to this disparity is the inability to do symbolic reasoning and program execution. GPT-4 lacks the capability to execute the given code to retrieve an output, which could help it gain deeper understanding of the underlying bugs. To mitigate this gap, we employ external tools to execute programs and extract useful symbolic information. We then supply this information to GPT-4 for feedback generation. Our approach centers on leveraging two categories of symbolic data: failing test cases and fixed programs.
Input/output for a failing test case. To highlight the error in the buggy program $\mathcal{P}_{\textnormal{b}}$, we provide GPT-4 with a test case for which $\mathcal{P}_{\textnormal{b}}$ fails to produce the expected output. To acquire this test case, we run $\mathcal{P}_{\textnormal{b}}$ on the existing test suite given for the corresponding task $\mathcal{T}$. The first test case in which $\mathcal{P}_{\textnormal{b}}$ fails is selected. We denote the triplet comprising this input, the output generated by $\mathcal{P}_{\textnormal{b}}$, and the expected output, as $\omega$ and include it in the prompt for feedback generation.
Fixed program. The fixed program, denoted as $\mathcal{P}_{\textnormal{f}}$, is generated using GPT-4, employing a procedure adapted from the work in [10]. To be more specific, we initiate the process by requesting the model to produce $10$ independent fixed programs. For this purpose, we include $\mathcal{T}$ and $\mathcal{P}_{\textnormal{b}}$ in the prompt [^3] The prompt used here has the same format as shown in Figure 4 (third prompt). to ask for 10 outputs (each output contains a fixed program) with the hyperparameter temperature set to $0.5$. Then, from this set of $10$, we take the programs that pass the test suite for $\mathcal{T}$ and among them, identify $\mathcal{P}_{\textnormal{f}}$ as the one with the smallest token-edit distance w.r.t. $\mathcal{P}_{\textnormal{b}}$. To compute the token-edit distance between two programs, we first tokenize them using Pygments library [31] and then calculate the Levenshtein edit distance based on the tokenized strings. If $\mathcal{P}_{\textnormal{f}}$ is found, we include it in the prompt for feedback generation. If, however, none of the generated programs is correct, we opt to exclude this symbolic information from the prompt.
@@ 3.2 Stage-2: Generate Feedback
Overview and intuition. In this stage, we aim to obtain a human tutor-style hint $\mathcal{H}$ as feedback to be given to the student, as previously mentioned in Section 2. In addition to our request for a hint $\mathcal{H}$ from GPT-4, we also ask for a detailed explanation, denoted as $\mathcal{X}$, for the bugs in $\mathcal{P}_{\textnormal{b}}$. The reason to ask for this explanation draws inspiration from Chain-of-Thought [18], an established method renowned for enhancing the reasoning capabilities of LLMs. The essence of the Chain-of-Thought approach lies in encouraging LLMs to explain their thought process meticulously, step by step, prior to presenting the final output. Within the specific context of hint generation, we allow the model to elaborate its reasoning through $\mathcal{X}$ before coming up with the concise single-sentence hint $\mathcal{H}$, which is essentially an abstracted version of the explanation. Furthermore, $\mathcal{X}$ will also play a pivotal role in the subsequent feedback validation stage, which will be elaborated upon in Section 3.3.
Prompt for feedback generation. In Figure 4 (first prompt), we provide our prompt for generating feedback. This prompt comprises the problem description for $\mathcal{T}$, the buggy program $\mathcal{P}_{\textnormal{b}}$, the symbolic information as extracted from the previous stage, and a request for an explanation $\mathcal{X}$ along with a hint $\mathcal{H}$. To get a response from GPT-4, we use this prompt while configuring the hyperparameter temperature to 00, indicating our preference for the most probable answer. All other hyperparameters are kept at their default settings. Following this, $\mathcal{X}$ and $\mathcal{H}$ are then extracted automatically from the output.